OpenStudy (big_mensa961):

it just says that I should expand that, to find the constant term. (2x-3/x^2)^9

Directrix (directrix):

Would you check to see if you posted the problem correctly? Thanks.

OpenStudy (big_mensa961):

where at plz

jhonyy9 (jhonyy9):

would it being (2x-3)^9 ((2x -3)^3)^3 ---------- = ----------------- = (x^2)^9 (x^2)^9 - in the above case use formula of (a-b)^3 = ? - this constant term will be this integer of -3 on exponent what will get at the end - i think @Directrix your opinion please about this ? thank you in advance

OpenStudy (big_mensa961):

Thanks you guys

OpenStudy (anonymous):

need help with another 1 mate

OpenStudy (big_mensa961):

not right now